2 people live 225 miles apart. The meet for lunch at a place between their houses. They leave at the same time. One takes 1hour and 45minutes to arrive and the other takes 1hour and 15 minutes to arrive. If they drive at the same speed determine their speed. And how far the resturant is from each house.
Answer by Paul(988)

1 h and 45 minutes = 1.75h.
1 h and 15 minutes = 1.25h.

1.75x+1.25x=225
3x=225
x=75
1.75*(75)=131.25
1.25*(75)=93.75

Hence there speeds are 75mph, and for the person whose time is 1h and 45 minutes his distance is 131.25miles and for the person whose time is 1.25 his distance is 93.75 miles.
PAul.